# Empowering Source-Free Domain Adaptation via MLLM-Guided Reliability-Based Curriculum Learning

### v2

@@ 4 Reliability-based Curriculum Learning
While pseudo-labels from MLLMs provide a strong starting point, their reliability varies across samples. Directly training on all pseudo-labels can introduce noise and hinder adaptation. To address this, we propose Reliability-based Curriculum Learning (RCL), shown in Figure 2, which builds on consensus-based reliability to strategically utilize target domain data. As shown in Figure 2, RCL consists of three stages: (1) Reliable Knowledge Transfer (RKT), (2) Self-correcting and MLLM-guided Knowledge Expansion (SMKE), and (3) Multi-hot Masking Refinement (MMR). In RKT (Sec.4.1), we begin by training the target model on the reliable subset $\mathcal{D}_{\textit{R}}$. Next, in SMKE (Sec.4.2), we fine-tune the model using both the reliable and less reliable subsets $\mathcal{D}_{\textit{R}}$. Finally, in MMR (Sec. 4.3), we refine the model on the full target dataset $\mathcal{D}_{t}$, incorporating the proposed Multi-hot Masking and consistency regularization for the unreliable subset $\mathcal{D}_{\textit{UR}}$. This staged learning strategy enables the model to first learn from reliable samples before gradually incorporating noisier pseudo-labels, reducing errors from premature exposure to unreliable data. By progressively refining its training, the model’s learning trajectory is effectively regularized, preventing instability that could arise from training on all data at once.
@@ 4.1 Reliable Knowledge Transfer (RKT)
We begin by transferring the most reliable MLLM knowledge to the target model. RKT trains the target model using pseudo-labels from the reliable subset $\mathcal{D}_{\textit{R}}$ in a supervised manner. At this stage, the model exclusively relies on high-confidence MLLM pseudo-labels, as it has not yet learned from the target domain. The reliable subset $\mathcal{D}_{\textit{R}}$ consists of samples for which all MLLMs agree on the pseudo-label:
$\mathcal{D}_{\textit{R}}=\{(x^{i}_{r},y^{i}_{r})\mid\mathcal{R}(x^{i}_{r})=1\},$ (4)
where $x^{i}_{r}$ is the $i$-th sample in the reliable subset, $\tilde{y}^{i}_{r}$ is the corresponding pseudo-label agreed upon by all MLLMs, and $\mathcal{R}(x^{i}_{r})$ is the reliability measure defined in the previous section. The target model $f_{\theta_{t}}$ is trained using a supervised cross-entropy loss on the reliable subset $\mathcal{D}_{\textit{R}}$:
@@ 4.2 Self-correcting and MLLM-guided Knowledge Expansion (SMKE)
Following RKT, RCL integrates less reliable pseudo-labels to expand the target model’s knowledge. Since the target model has been pre-trained on the source model and fine-tuned with RKT, we transition from direct distillation to a more adaptive learning process. When the target model is confident, it refines its predictions through self-correction. In cases of lower confidence, MLLMs serve as guidance rather than fixed teachers, allowing the model to expand its knowledge dynamically.
To facilitate this, we propose SMKE, which fine-tunes the target model on both the reliable and less reliable subsets, $\mathcal{D}_{R}\cup\mathcal{D}_{LR}$. This enables the target model to learn from a larger portion of the target domain data and benefit from the additional information provided by $\mathcal{D}_{LR}$. The pseudo-label $\tilde{y}^{i}$ used for training the target model is determined based on the confidence of the target model’s predictions. Let $\hat{y}_{t}^{i}$ be the pseudo-label predicted by the target model for the target domain sample $x_{t}^{i}$, and let $p_{t}^{i}$ be the corresponding confidence score, which is calculated as the maximum value of the target model’s predictive probabilities. We define a confidence threshold $\tau$ to determine whether to use the target model’s pseudo-label or the MLLMs’ pseudo-label. The pseudo-label $\tilde{y}^{i}$ is as follows:
@@
$\mathcal{L}_{\text{SMKE}}=-\frac{1}{|\mathcal{D}_{\textit{R}}\cup\mathcal{D}_{\textit{LR}}|}\sum_{x^{i}_{t}\in\{\mathcal{D}_{\textit{R}}\cup\mathcal{D}_{\textit{LR}}\}}\tilde{y}^{i}\cdot\log f_{\theta_{t}}(x^{i}_{t}),$ (7)
@@ 4.3 Multi-hot Masking Refinement (MMR)
Finally, we expand to the full training set to incorporate the unreliable subset $\mathcal{D}_{UR}$, which consists of target domain samples where the MLLMs disagree on the pseudo-labels. This disagreement makes it challenging to assign reliable pseudo-labels to the samples in $\mathcal{D}_{\textit{UR}}$. Thus, these samples are not included during RKT and SMKE. To fully exploit $\mathcal{D}_{t}$, we propose Multi-hot Masking Refinement (MMR), integrating joint selection, multi-hot masking, and consistency regularization to utilize these challenging samples. The full algorithm for MMR is shown in Algo. 1.
@@
$\mathcal{L}_{\text{cons}}=\frac{1}{M}\sum_{m=1}^{M}\sum_{i=1}^{N_{t}}\text{H}(\tilde{y}^{i},\mathbf{z}_{st}^{i}),$ (10)

### v3

@@ 4 Reliability-based Curriculum Learning
Pseudo-labels from teachers provide a strong starting point, but their reliability can vary across samples. To address this, we propose Reliability-based Curriculum Learning (RCL) (Fig. 4), which uses consensus-based reliability to progressively incorporate target data by pseudo-label confidence. RCL consists of three stages: (1) Reliable Knowledge Transfer (RKT) trains on the most reliable subset $\mathcal{D}_{\textit{R}}$; (2) Self-correcting and MLLM-guided Knowledge Expansion (SMKE) integrates less reliable samples with adaptive correction; and (3) Multi-hot Masking Refinement (MMR) incorporates the full target set $\mathcal{D}_{t}$, using a union-of-classes mask with consistency regularization to handle the most uncertain subset $\mathcal{D}_{\textit{UR}}$. This staged curriculum allows the student to learn from confident samples first and gradually expand to noisier ones, stabilizing adaptation. RCL is teacher-agnostic, being a multi-teacher curriculum designed to handle noisy and unreliable pseudo-labels. This makes the framework broadly applicable to CLIP or other models. Our STS conversion integrates MLLMs by adapting their free-form outputs into usable class predictions. We view MLLMs as an upper-bound case: their foundation knowledge and demonstrated cross-domain generalization offer a strong motivation for SFDA in real-world settings such as medical applications.
@@ 4.1 Reliable Knowledge Transfer (RKT)
In the first stage, we train the student on the reliable subset $\mathcal{D}_{\textit{R}}$ where multiple teachers fully agree. This ensures that early learning is guided by high-quality labels and avoids propagating noise. The intuition is to give the student a “clean start”: by only learning from samples with consensus, the model builds a stable foundation before being exposed to more uncertain data. The reliable subset $\mathcal{D}_{\textit{R}}$ consists of samples for which all MLLMs agree on the pseudo-label $\mathcal{D}_{\textit{R}}=\{(x^{i}_{r},y^{i}_{r})\mid\mathcal{R}(x^{i}_{r})=1\},$ where $x^{i}_{r}$ is the $i$-th sample in the reliable subset, $\tilde{y}^{i}_{r}$ is the corresponding pseudo-label agreed upon by all MLLMs, and $\mathcal{R}(x^{i}_{r})\in[0,1]$ is the reliability measure defined in the previous section. The target model $f_{\theta_{t}}$ is trained using a supervised cross-entropy loss on the reliable subset $\mathcal{D}_{\textit{R}}$:
@@ 4.2 Self-correcting and MLLM-guided Knowledge Expansion (SMKE)
Following RKT, RCL expands training by incorporating less reliable pseudo-labels to broaden the student’s knowledge. Since the student has already been pre-trained and stabilized through RKT, we transition from strict teacher supervision to a more adaptive learning process. When the student is confident, it self-corrects its predictions; when confidence is low, MLLMs provide auxiliary guidance rather than acting as fixed teachers. The intuition is that after RKT, the student has matured enough to act as a co-teacher: its predictions can be trusted to correct or supplement partial teacher agreement, enabling a gradual and safer expansion of usable data. To implement this, SMKE fine-tunes the model on both reliable and less reliable subsets, $\mathcal{D}_{R}\cup\mathcal{D}_{LR}$. For each sample $x_{t}^{i}$, the pseudo-label $\tilde{y}^{i}$ is determined adaptively: if the student’s prediction $\hat{y}_{t}^{i}$ has confidence $p_{t}^{i}$ above a threshold $\tau$, we use it; otherwise, we defer to the MLLM teachers’ pseudo-label. This balances the student’s growing reliability with external guidance, reducing dependence on any single source. The pseudo-label $\tilde{y}^{i}$ is as follows:
@@
$\mathcal{L}_{\text{SMKE}}=-\frac{1}{|\mathcal{D}_{\textit{R}}\cup\mathcal{D}_{\textit{LR}}|}\sum_{x^{i}_{t}\in\{\mathcal{D}_{\textit{R}}\cup\mathcal{D}_{\textit{LR}}\}}\tilde{y}^{i}\cdot\log f_{\theta_{t}}(x^{i}_{t}).$ (5)
@@ 4.3 Multi-hot Masking Refinement (MMR)
Finally, we refine the student on the full target dataset $\mathcal{D}_{t}$, focusing on the most uncertain subset $\mathcal{D}_{\textit{UR}}$. Naively minimizing entropy on these samples risks overfitting to noise or distractors. MMR instead narrows the prediction space by masking logits to the union of teacher-suggested classes, forcing the model to choose only among classes deemed plausible by at least one teacher. We further apply consistency regularization between weak and strong augmentations to stabilize training when agreement is low. The key intuition is that by this stage the student has already learned reliable patterns (via RKT and SMKE), so its own predictions are valuable. MMR leverages both the student’s maturity and the teachers’ broad guidance: disagreement is treated as ambiguity rather than error, and the mask provides a constrained search space for refinement. This enables the model to still benefit from hard samples while reducing error propagation.
Table 1: Accuracy (%) on Office-Home. SF: source-free; CP, ViT: CLIP, ViT backbone. Best results are bolded, second-best underlined. (*) indicates zero-shot performance. Full table in Supplementary Material, Tab. 7.
[…]
Source - 44.7 64.2 69.4 48.3 57.9 60.3 49.5 40.3 67.2 59.7 45.6 73.0 56.7
PADCLIP-RN [lai2023padclip] 57.5 84.0 83.8 77.8 85.5 84.7 76.3 59.2 85.4 78.1 60.2 86.7 76.6
ADCLIP-RN [singha2023ad] 55.4 85.2 85.6 76.1 85.8 86.2 76.7 56.1 85.4 76.8 56.1 85.5 75.9
ELR [yi2023source] 58.4 78.7 81.5 69.2 79.5 79.3 66.3 58.0 82.6 73.4 59.8 85.1 72.6
PLUE [litrico_2023_CVPR] 49.1 73.5 78.2 62.9 73.5 74.5 62.2 48.3 78.6 68.6 51.8 81.5 66.9
C-SFDA [karim2023c] 60.3 80.2 82.9 69.3 80.1 78.8 67.3 58.1 83.4 73.6 61.3 86.3 73.5
PSAT-GDA [tang2023progressive] 73.1 88.1 89.2 82.1 88.8 88.9 83.0 72.0 89.6 83.3 73.7 91.3 83.6
DIFO-C-RN [tang2023sourceMLLM] 62.6 87.5 87.1 79.5 87.9 87.4 78.3 63.4 88.1 80.0 63.3 87.7 79.4
DIFO-C-B32 [tang2023sourceMLLM] 70.6 90.6 88.8 82.5 90.6 88.8 80.9 70.1 88.9 83.4 70.5 91.2 83.1
CLIP-RN [CLIP_2021]* - 51.7 85.0 83.7 69.3 85.0 83.7 69.3 51.7 83.7 69.3 51.7 85.0 72.4
LLaVA-34B (w/ STS) [liu2023llava]* - 78.3 93.7 89.5 87.0 93.7 89.5 87.0 78.3 89.5 87.0 78.3 93.7 87.2
InstBLIP-XXL (w/ STS) [dai2024instructblip]* - 82.0 91.6 88.8 82.2 91.6 88.8 82.2 82.0 88.8 82.2 82.0 91.6 86.2
ShrGPT4V-13B (w/ STS) [chen2023sharegpt4v]* - 66.7 85.8 84.8 83.2 85.8 84.8 83.2 66.7 84.8 83.2 66.7 85.8 80.1
Table 2: Accuracy (%) on DomainNet and average accuracy on VisDA. Full table in Supplementary Material, Tab. 8,  9.
Method SF CP ViT DomainNet VisDA
[…]
Source - 42.6 53.7 51.9 52.9 66.7 51.6 49.1 56.8 43.9 60.9 48.6 53.2 52.7 45.3
DAPL-RN [ge2023domain] 72.4 87.6 65.9 72.7 87.6 65.6 73.2 72.4 66.2 73.8 72.9 87.8 74.8 86.9
ADCLIP-RN [lai2023padclip] 71.7 88.1 66.0 73.2 86.9 65.2 73.6 73.0 68.4 72.3 74.2 89.3 75.2 88.5
PLUE [litrico_2023_CVPR] 59.8 74.0 56.0 61.6 78.5 57.9 61.6 65.9 53.8 67.5 64.3 76.0 64.7 88.3
TPDS [tang2023source] 62.9 77.1 59.8 65.6 79.0 61.5 66.4 67.0 58.2 68.6 64.3 75.3 67.1 87.6
DIFO-C-RN [tang2023sourceMLLM] 73.8 89.0 69.4 74.0 88.7 70.1 74.8 74.6 69.6 74.7 74.3 88.0 76.7 88.8
DIFO-C-B32 [tang2023sourceMLLM] 76.6 87.2 74.9 80.0 87.4 75.6 80.8 77.3 75.5 80.5 76.7 87.3 80.0 90.3
LLaVA-34B (w/ STS) [liu2023llava]* - 84.4 91.0 83.7 85.5 91.0 83.7 85.5 84.4 83.7 85.5 84.4 91.0 86.1 92.1
InstBLIP-XXL (w/ STS) [dai2024instructblip]* - 82.5 89.0 83.0 86.7 89.0 83.0 86.7 82.5 83.0 86.7 82.5 89.0 85.3 86.7
ShrGPT4V-13B (w/ STS) [chen2023sharegpt4v]* - 79.7 87.9 79.2 79.9 87.9 79.2 79.9 79.7 79.2 79.9 79.7 87.9 81.7 90.4
RCL (Ours) 87.6 92.8 87.9 89.2 92.7 87.8 89.6 87.7 87.6 89.4 87.5 92.7 89.4 93.2
RCL-ViT (Ours) 88.1 93.3 88.0 89.7 93.3 88.0 89.7 88.0 87.8 89.7 88.1 93.3 89.7 93.3
@@
$\mathcal{L}_{\text{cons}}=\frac{1}{M}\sum_{m=1}^{M}\sum_{i=1}^{N_{t}}\mathcal{L}_{\text{CE}}(\tilde{y}^{i},\mathbf{z}_{st}^{i}),$ (8)